### v3

# Eigenvalue and pseudospectrum processes generated by nonnormal Toeplitz matrices with rank 1 perturbations

Abstract
Motivated by recent studies on a time-dependent random matrix model called the non-Hermitian matrix-valued Brownian motion, we propose two kinds of dynamical processes of $n\times n$ matrices generated by nonnormal Toeplitz matrices. As a deterministic rank 1 perturbation, $\delta J$ is added at time $m\leq n$, where $\delta\in{\mathbb{C}}$ and $J$ is the all-ones matrix. For each process, first we report the complicated motion of the eigenvalues which are numerically calculated. Then we derive the specific equation which determines the motion of non-zero eigenvalues and clarifies the time-dependence of degeneracy of the zero-eigenvalue. The Catalan numbers play an important role to describe an outlier of non-zero eigenvalues. Comparison with the solutions of this equation, it is concluded that the numerically observed non-zero eigenvalues distributing around the zero-eigenvalue are not the exact eigenvalues of the original system but of the system perturbed by uncontrolled rounding errors of computer. This phenomenon implies the high sensitivity of the degenerated zero-eigenvalue to random dense perturbations. The domain in which the eigenvalues of randomly perturbed system are distributed is identified with the pseudospectrum including the zero-eigenvalue of the original system with $\delta J$. We characterize the pseudospectrum processes using the symbol curves of the corresponding nonnormal Toeplitz operators without $\delta J$. We report new phenomena in our second model such that at each time the outermost closed simple curve cut out from the symbol curve is realized as the exact eigenvalues, but the inner part of symbol curve is reduced in size and embedded in the pseudospectrum including the zero-eigenvalue. The asymptotics of the processes in the infinite-matrix limit $n\to\infty$ are also discussed.
Keywords: Nonnormal Toeplitz matrices, Eigenvalue processes, Pseudospectrum processes, Symbol curves of Toeplitz operators, Catalan numbers
1 Introduction
The present work was motivated by recent studies on a time-dependent random matrix model called the non-Hermitian matrix-valued Brownian motion (BM) [12, 13, 14, 15, 18, 21, 29]. Let $n\in\mathbb{N}:=\{1,2,\dots\}$. Consider $2n^{2}$ independent one-dimensional standard BMs, $(B^{{\rm R}}_{jk}(t))_{t\geq 0}$, $(B^{{\rm I}}_{jk}(t))_{t\geq 0}$, $1\leq j,k\leq n$. Let $i:=\sqrt{-1}$ and we define the $n\times n$ non-Hermitian matrix-valued BM by
$M(t)=(M_{jk}(t))_{1\leq j,k\leq n}:=\left(\frac{1}{\sqrt{2n}}(B^{{\rm R}}_{jk}(t)+iB^{{\rm I}}_{jk}(t))\right)_{1\leq j,k\leq n},\quad t\geq 0,$
which starts from a deterministic matrix, $M(0)=(M_{jk}(0))_{1\leq j,k\leq n}\in\mathbb{C}^{n^{2}}$. We consider the eigenvalue process of $(M(t))_{t\geq 0}$ denoted by $\bm{\Lambda}(t)=(\Lambda_{j}(t))_{j=1}^{n}\in\mathbb{C}^{n}$, $t\geq 0$. When the matrix-valued BM starts from the null matrix, $M(0)=O$, $(\bm{\Lambda}(t))_{t\geq 0}$ starts from the $n$ particles all degenerated at the origin, $n\delta_{0}$, and exhibits a uniform distribution in an expanding disk centered at the origin with radius $\sqrt{t}$ on a complex plane $\mathbb{C}$, $t>0$ (the circular law) [4, 27, 38]. At each time $t>0$, $\Lambda(t)$ is identified with the complex Ginibre ensemble of eigenvalues with variance $t$ [26], which has been extensively studied in random matrix theory (see [16, 19] and the references therein).
Burda et al. [15] studied the process starting from
$M(0)=S,$
where $S$ denotes the $n\times n$ shift matrix
[…]
where $\delta_{jk}$ denotes the Kronecker delta. Notice that this matrix is nonnormal; $S^{\dagger}S\not=SS^{\dagger}$. (In the present paper, for a square matrix $A\in\mathbb{C}^{n^{2}}$, $A^{\dagger}$ denotes the complex conjugate of transpose of $A$; $A^{\dagger}:=\overline{A^{\sf T}}$.) All eigenvalues of $S$ are zero, and hence the initial state of $(\bm{\Lambda}(t))_{t\geq 0}$ is $n\delta_{0}$, which is the same as that in the case $M(0)=O$. By numerical simulation, however, Burda et al. [15] found that the eigenvalues seem to expand instantly from $n\delta_{0}$ to make a unit circle as dotted in Fig. 1. For the time interval $10<t<$ 1, the dots form the growing annulus (Fig. 1). Then the inner radius of the annulus shrinks to zero at $t=1$ and dots fill up a full disk (Fig. 1). The disk expands with radius $\sqrt{t}$, $t>1$ and the dots tend to follow the circular law in $t\gg 1$.
(a) Ring structure ($0<t\ll 1$)
(b) Growing annulus ($10<t<$ 1)
(c) Fulfilled disk ($t\geq 1$)
Figure 1: Numerically obtained eigenvalues are dotted for the non-Hermitian matrix-valued Brownian motion, $(M(t))_{t\geq 0}$, with size $n=300$ starting from the nonnormal matrix $S$.
It was claimed that [15] the unit circle observed at the early stage will be composed not of exact eigenvalues but of pseudoeigenvalues [11, 34, 39]. For a complex-valued square matrix $A\in\mathbb{C}^{n^{2}}$ with an arbitrary but fixed $\varepsilon>0$, the $\varepsilon$-pseudospectrum of $A$ is defined as an open subset $\sigma_{\varepsilon}(A)$ of $z\in\mathbb{C}$ such that
$\|(zI-A)^{-1}\|>\varepsilon^{-1}.$ (1.2)
[…]
$\|A\|:=\max_{\bm{x}\in\mathbb{C}^{n}}\frac{\|A\bm{x}\|_{2}}{\|\bm{x}\|_{2}},$
where $\|\bm{x}\|_{2}:=\sqrt{\sum_{j=1}^{n}|x_{j}|^{2}}$ for vector $\bm{x}=(x_{1},\dots,x_{n})\in\mathbb{C}^{n}$. If $\sigma(A)$ denotes the spectra (i.e. set of eigenvalues of $A$), then
$\|(zI-A)^{-1}\|=\infty,\quad z\in\sigma(A).$
Hence, by definition, the exact eigenvalue is contained in the $\varepsilon$-pseudospectrum for every $\varepsilon>0$. In other words, the eigenvalues of $A$ are recovered from $\sigma_{\varepsilon}(A)$ as poles of the resolvent $(z-A)^{-1}$ in the $\varepsilon\to 0$ limit [39].
It is proved that the above definition of $\varepsilon$-pseudospectra with (1.2) is equivalent with the following definition [34] [39, Theorem 2.1]. The $\varepsilon$-pseudospectrum of a matrix $A\in\mathbb{C}^{n^{2}}$ with $\varepsilon>0$ is the set of $z\in\mathbb{C}$ such that
$z\in\sigma(A+E)$
for some matrix $E\in\mathbb{C}^{n^{2}}$ with $\|E\|<\varepsilon$. That is, for a given matrix, the $\varepsilon$-pseudospectrum is not the exact spectrum of the original matrix $A$, but it is the set of exact eigenvalues of some perturbed matrix $A+E$ with $\|E\|<\varepsilon$. We notice that, at $0<t\ll 1$, the matrix $M(t)$ starting from (1.1) will be well approximated by
$S+\sqrt{\frac{t}{2n}}Z,$ (1.3)
where $Z=(Z_{jk})_{1\leq j,k\leq n}$ with the elements given by independent complex Gaussian random variables,
$Z_{jk}=X_{jk}+iY_{jk},\quad X_{jk}\sim{\rm N}(0,1),\quad Y_{jk}\sim{\rm N}(0,1),\quad 1\leq j,k\leq n.$
The equivalence of the two definitions of pseudospectrum mentioned above suggests the following: The eigenvalues with the ring structure reported by Burda et al. [15] will be the eigenvalues of the randomly perturbed matrix (1.3) and the domain surrounded by the ring is the pseudospectrum including the highly degenerated zero-eigenvalue of $M(0)=S$ [32]. As a generalization of the randomly perturbed system (1.3), we are interested in the discrete-time random process,
$S_{\delta Z}(m):=S^{m}+\delta Z,\quad m=1,2,\dots,n,$ (1.4)
where the nonnormality of the matrix is changing in time $m$.
In the present paper, we introduce the following two kinds of discrete-time dynamical systems generated by nilpotent Toeplitz matrices, in which the random perturbations by $Z$ are replaced by the deterministic rank 1 perturbations [20],
[…]
$:=S^{m}+aS^{m+1}+\delta J,\quad m=1,2,\dots,n,$
where $\delta,a\in\mathbb{C}$ and $J=(J_{jk})_{1\leq j,k\leq n}$ is the all-ones matrix;
$J_{jk}\equiv 1,\quad 1\leq j,k\leq n.$
The model 1 is a deterministic version of (1.4) and model 2 is its one-parameter ($a\in\mathbb{C}$) extension. Notice that $S^{(\ell)}_{\delta J}(m)$, $\ell=1,2$ are non-Hermitian; $(S^{(\ell)}_{\delta J}(m))^{\dagger}\not=S^{(\ell)}_{\delta J}(m)$, and nonnormal,
$(S^{(\ell)}_{\delta J}(m))^{\dagger}S^{(\ell)}_{\delta J}(m)\not=S^{(\ell)}_{\delta J}(m)(S^{(\ell)}_{\delta J}(m))^{\dagger},\quad m=1,2,\dots,n-1.$
For nonnormal matrices, both of right-eigenvectors and left-eigenvectors are needed to construct eigenspaces associated with eigenvalues, and the overlap matrix is defined using the right- and left-eigenvectors. The overlap matrices play important roles in a variety of fields in mathematics and physics. See [2, 10, 12, 13, 14, 15, 16, 17, 21, 22, 23, 24, 25, 29, 30, 31, 40] and references therein. In particular, in the recent study of non-Hermitian matrix-valued stochastic processes, the analysis of the coupling between the eigenvalue processes and the eigenvalue-overlap processes is one of the central topics [12, 13, 14, 15, 18, 21, 29]. The square roots of the diagonal elements of the overlap matrix are especially called the condition numbers of eigenvalues and we notice that the pseudospectra can be evaluated using the condition numbers (e.g. the Bauer–Fike theorem [39, Sects. 35 and 52]). In the present paper, we study the coupling between the eigenvalue processes and the pseudospectrum processes for the two models.
For $m=2,3,\dots,n$, both models have a degenerated eigenvalue $\lambda_{0}=0$. There are two notions of degeneracy of eigenvalues. The algebraic multiplicity is the number of times the eigenvalue appears as a root of the characteristic polynomial of the matrix, while the geometric multiplicity is the dimension of the linear space of the eigenvectors associated to the eigenvalue. If the algebraic multiplicity of an eigenvalue exceeds the geometric multiplicity, then that eigenvalue is said to be defective and the matrix becomes nondiagoralizable [39]. In both models, we can prove that $\lambda_{0}$ is defective if $m=2,3,\dots,n-2$ [32]. The defectivity of $\lambda_{0}$ is expressed by the size of $\varepsilon$-pseudospectrum including $\lambda_{0}$ for given $0<\varepsilon\ll 1$.
If $\delta=0$, that is, the deterministic rank 1 perturbation $\delta J$ is not applied, then the matrices of our models are reduced to banded Toeplitz matrices for any $n\in\mathbb{N}$ and their $n\to\infty$ limits define the banded Toeplitz operators. For any banded Toeplitz operator, a closed curve called the symbol curve is defined on the complex plane $\mathbb{C}$ and the $\varepsilon$-pseudospectra of the nonnormal banded Toeplitz matrices are specified by the whole regions surrounded by the symbol curves [34, 39]. In the present paper we report new phenomena in model 2 with $\delta\not=0$ such that at each time $m=2,3,\dots n-2$, the outermost closed simple curve cut out from the symbol curve is realized as the exact eigenvalues, but the inner part of symbol curve composed of several closed simple curves osculating each other is reduced in size and embedded as a complicated structure in the pseudospectrum including $\lambda_{0}$. The size reduction of the pseudospectrum including $\lambda_{0}$ expresses the relaxation process of defectivity of $\lambda_{0}$. Our dynamical systems show the transitions from far-from-normal matrices to near-normal matrices. Algebraic descriptions of such relaxation processes of defectivity is studied in [32] by calculating the Jordan block decompositions of the resolvents of our matrices at $\lambda_{0}$. We expect that the present mathematical study of eigenvalue and pseudospectrum processes will provide useful models for physical phenomena studied in non-Hermitian quantum physics [3, 28, 33].
The paper is organized as follows. In Section 2 we report the numerical observations for the eigenvalue processes for the models. In Section 3 we give the equations which determine the exact eigenvalues in Theorem 3.2 for model 1 and in Theorem 3.5 for model 2, respectively. The properties of the solutions of these equations are given by Propositions 3.3, 3.4, and 3.7. The notions of symbols and symbol curves for Toeplitz operators are introduced in Section 4 and our numerical results are studied in referring to the theory of pseudospectra for banded nonnormal Toeplitz matrices. There we study new phenomena in pseudospectrum processes exhibiting separation of symbol curves and dilatation of their inner parts. They have not been reported in the previous mathematical study of banded-Toeplitz-matrix models with random perturbations [5, 6, 7, 8, 9, 36]. Section 5 is devoted to reporting the asymptotics of the exact-eigenvalue processes and pseudospectrum processes in infinite-matrix limit $n\to\infty$ (Propositions 5.1 and 5.2). Concluding remarks and future problems are given in Section 6, where a possibility to extend the present study to the random process (1.4) is discussed.
2 Numerical Observations of Processes
2.1 Model 1
We have performed numerical calculation of the eigenvalue processes of model 1 with a given size of matrix $n\in\mathbb{N}$. The obtained eigenvalues are plotted on $\mathbb{C}$ for each time $m=1,2,\dots,n$.
Figure 2: Plots of numerically obtained eigenvalues at time $m=1$ for model 1, $S^{(1)}_{\delta J}(1)$, with $n=200$ and $\delta=0.01$. 199 dots form a unit circle missing one point at $z=1$, and an outlier is observed near $z=3$, which is denoted by $\lambda_{1}(1)$.
(a) $m=2$
(b) $m=8$
(c) $m=15$
(d) $m=80$
Figure 3: Numerically obtained eigenvalues are plotted for model 1, $(S^{(1)}_{\delta J}(m))_{m=1}^{n}$, with $n=200$ and $\delta=0.01$ at $m=2$, 8, 15, and 80, respectively.
The observations are explained using the case with $n=200$ and $\delta=0.01$ below.
(i)  The numerically obtained eigenvalues are dotted in Fig. 2 at $m=1$. We find 199 dots which form a unit circle missing one point at $z=1$, and one outlier located near $z=3$.
As proved by Propositions 3.3 (iia) and 3.4 in Section 3, the dot near $z=3$ is identified with an exact eigenvalue of $S_{\delta J}^{(1)}(1)$ and its time evolution, denoted by $(\lambda_{1}(m))_{m=1}^{n}$, can be explicitly described using the Catalan numbers ${C_{k}:=\frac{1}{k+1}\binom{2k}{k}=\frac{(2k)!}{(k+1)!k!}}$, $k\in\mathbb{N}_{0}$ [35] as
$\lambda_{1}(m)$ $=n\delta+1-\sum_{k=0}^{p_{1}-1}C_{k}\left(\frac{m}{n}\right)^{k+1}\frac{1}{(n\delta)^{k}}+{\rm O}((n\delta)^{-p_{1}}),$
for $11\leq m\leq n-$ 1, where $p_{1}:=[(n-1)/m]$ (the greatest integer less than or equal to $(n-1)/m$), and $\lambda_{1}(n)=n\delta$ (Proposition 3.4). So we will show mainly the eigenvalues which are distinct from $\lambda_{1}(m)$ in the following figures. (For outliers in spectra discussed in random matrix theory, see [20, 37].)
Figure 3 shows the numerical results for $m\geq 2$.
(ii)   At time $m=2$, 99 dots form a slightly deformed circle whose radius is $\lesssim 1$. In addition to them many dots appear and form a smaller circle with radius $\simeq 0.7$. In the vicinity of the origin, three dots are observed, one of which is located exactly at the origin.
(iii)   At time $m=8$, 24 dots form an incomplete circle shaped ‘C’, whose radius is about 0.8. In addition, we see a small annulus whose boundaries are wavy. Within that smaller circle, we see three dots, one of which is located at the origin.
(iv)   As $m$ increases, the inner annulus shrinks to the origin. The reduction of size is exponential as a function of $m$. The outer dots in the upper (resp. lower) half plane of $\mathbb{C}$ move along a circle counterclockwise (resp. clockwise) until they attach the negative real axis, :=$\mathbb{R}_{-}:=\{x\in\mathbb{R};x<0\}$. Then they move along $\mathbb{R}_{-}$ repulsively with each other preserving the order of the distances from the origin. They are absorbed by the origin one by one. At time $m=15$, only 13 dots remain apart from the origin. One of them is on $\mathbb{R}_{-}$, which will approach the origin and will be absorbed earlier than other 12 dots.
(v)   At time $m=80$, only two dots remain apart from the dot at the origin. In $m>80$, both of the two dots approach to $\mathbb{R}_{-}$, and then they show repulsive motion on $\mathbb{R}_{-}$. One of them is absorbed by the origin at $m=100$.
(vi)   At the final time $m=n=200$, there are only two dots, one of them is at the origin, and the other one is $\lambda_{1}(n)$ located at $z=2$. They are the eigenvalues of $\delta J$, since $S^{n}=0$ for the $n\times n$ shift matrix (1.1).
2.2 Model 2
(a) $m=1$
(b) $m=2$
(c) $m=3$
(d) $m=4$
Figure 4: Numerically obtained eigenvalues are plotted for model 2, $(S^{(2)}_{\delta J}(m))_{m=1}^{n}$, with $n=200$, $\delta=0.01$, and $a=1$ at time $m=1$, 2, 3, and 4, respectively.
In Fig. 4, we dotted the numerically obtained eigenvalues of model 2 for $m=1,2,3$, and 4. At $m=1$, a limaçon-like curve [34, 39] is observed. At $m=2,3$, and 4, a deformed circle whose radius is slightly less than 2 is formed by dots whose number decreases as $m$ increases. Notice that an outlier eigenvalue exists near $z=4$ in all figures. The structure found in the vicinity of the origin becomes more complicated as $m$ increases. This inner structure shrinks rapidly to the origin when $m\geq 5$. The motion of the outer dots is very similar to that observed in model 1: They move along the upper- or lower-half deformed circles to $\mathbb{R}_{-}$, show repulsive motion on $\mathbb{R}_{-}$, and then they are absorbed by the origin one by one.
3 Exact Eigenvalue Processes
3.1 Model 1
For model 1, we consider the following eigenvalue-eigenvector equations,
$S_{\delta J}(m)\bm{v}(m)=\lambda(m)\bm{v}(m),\quad m=1,2,\dots,n.$ (3.1)
Let ${\bf 1}$ be the all-ones vector and we introduce the Hermitian inner produce, $\langle\bm{u},\bm{v}\rangle:=\sum_{j=1}^{n}u_{j}\overline{v_{j}}$, $\bm{u},\bm{v}\in\mathbb{C}^{n}$. Define
$\alpha(m):=\langle\bm{v}(m),{\bf 1}\rangle=\sum_{j=1}^{n}v_{j}(m).$
Let $1_{(\omega)}$ be the indicator function of the condition $\omega$; $11_{(\omega)}=$ 1 if $\omega$ is satisfied, and $01_{(\omega)}=$ 0 otherwise. The following fact will be used.
Lemma 3.1
For $\ell\in\mathbb{N}$,
$\langle S^{\ell}{\bf 1},{\bf 1}\rangle=(n-\ell)1_{(1\leq\ell\leq n-1)}.$
For $m\in\{1,2,\dots,n\}$, define
$\widehat{p}(n,m):=\frac{n}{m}\quad\mbox{and}\quad p(n,m):=[\widehat{p}(n,m)],$ (3.2)
where $[x]$ denotes the greatest integer less than or equal to $x\in\mathbb{R}$ (the floor function of $x$). Let
$p_{1}:=p(n-1,m).$ (3.3)
The following is proved.
Theorem 3.2
For $m\in\{1,2,\dots,n\}$, there are $p_{1}+1$ non-zero eigenvalues, which solve the following equation,
$\frac{1}{n\delta}z^{p_{1}+1}-\frac{1-z^{p_{1}+1}}{1-z}+\frac{m}{n}\frac{1}{1-z}\left\{p_{1}+1-\frac{1-z^{p_{1}+1}}{1-z}\right\}=0.$ (3.4)
This equation is also written as the polynomial equation,
$z^{p_{1}+1}-n\delta\sum_{k=0}^{p_{1}}\left\{1-(p_{1}-k)\frac{m}{n}\right\}z^{k}=0.$ (3.5)
The corresponding eigenvectors satisfy $\alpha(m)\not=0$. All other $n-p_{1}-1$ eigenvalues degenerate at zero. In this case, the corresponding eigenvectors satisfy $\alpha(m)=0$, that is, they are orthogonal to ${\bf 1}$.
Proof   Equation (3.1) is written as
$(zI-S^{m})\bm{v}(m)=\delta\alpha(m){\bf 1},$ (3.6)
for $z=\lambda(m)$, where we noticed the equality,
$J\bm{v}(m)=\alpha(m){\bf 1}.$ (3.7)
If we consider the zero-eigenvalue $z=\lambda(m)=0$, (3.6) becomes $-S^{m}\bm{v}(m)=\delta\alpha(m){\bf 1}$. Since $S^{m}$ shifts the elements of any vector upward by $m$ when $S^{m}$ is operated on the vector from the left, the last $m$ elements of the vector $-S^{m}\bm{v}(m)$ are zero. Since ${\bf 1}$ is the all-ones vector, $\alpha(m)$ should be 0. For non-zero eigenvalues $z=\lambda(m)\not=0$ on the other hand, we can assume $\alpha(m)\not=0$. We solve this equation as follows,
$\bm{v}(m)=\delta\alpha(m)(zI-S^{m})^{-1}{\bf 1}=\delta\alpha(m)\sum_{k=0}^{\infty}z^{-(k+1)}S^{mk}{\bf 1},$
where we used the expansion formula of an inverse matrix. By taking inner products with ${\bf 1}$ on both sides, we have
$\langle\bm{v}(m),{\bf 1}\rangle=\delta\alpha(m)\sum_{k=0}^{\infty}z^{-(k+1)}\langle S^{mk}{\bf 1},{\bf 1}\rangle.$
Since $mk\leq n-1,m,k\in\mathbb{N}\iff k\leq p_{1},k\in\mathbb{N},$ where $p_{1}$ is defined by (3.3), Lemma 3.1 gives
$\alpha(m)=\delta\alpha(m)\sum_{k=0}^{p_{1}}z^{-(k+1)}(n-mk).$ (3.8)
Here the fact $\langle\bm{v}(m),{\bf 1}\rangle=\alpha(m)$ was used. For non-zero eigenvalues, we can assume $\alpha(m)\neq 0$. Hence we have
$0\displaystyle 1=\delta\sum_{k=0}^{p_{1}}z^{-(k+1)}(n-mk)\iff\frac{1}{n\delta}-z^{-1}\sum_{k=0}^{p_{1}}z^{-k}+\frac{m}{n}z^{-1}\sum_{k=0}^{p_{1}}kz^{-k}=$ 0 . (3.9)
Then we use the summation formulas,
$z^{-1}\sum_{k=0}^{p_{1}}z^{-k}$ $=z^{-1}\frac{1-z^{-(p_{1}+1)}}{1-z^{-1}}=z^{-(p_{1}+1)}\frac{1-z^{p_{1}+1}}{1-z},$
$z^{-1}\sum_{k=0}^{p_{1}}kz^{-k}$ $=-\frac{\partial}{\partial z}\sum_{k=0}^{p_{1}}z^{-k}=\frac{z^{-(p_{1}+1)}}{1-z}\left\{p_{1}+1-\frac{1-z^{p_{1}+1}}{1-z}\right\}.$ (3.10)
Thus (3.9) is written as
$\frac{1}{n\delta}-z^{-(p_{1}+1)}\frac{1-z^{p_{1}+1}}{1-z}+\frac{m}{n}\frac{z^{-(p_{1}+1)}}{1-z}\left\{p_{1}+1-\frac{1-z^{p_{1}+1}}{1-z}\right\}=0.$
By multiplying $z^{p_{1}+1}$, (3.4) is obtained. Using the formulas (3.10), (3.5) is readily derived from (3.4). ∎
Remark 3.1   Fix $m\in\{1,2,\dots,n\}$. Suppose $\alpha(m)=0$, that is, $\bm{v}(m)$ is orthogonal to ${\bf 1}$. By (3.7), the eigenvalue-eigenvector equation (3.1) for $S_{\delta J}$ is reduced to that for $S^{m}$,
$S^{m}\bm{v}(m)=\lambda(m)\bm{v}(m).$
Consider the vectors in the form
$\bm{v}^{0}=(v^{0}_{1},v^{0}_{2},\cdots,v^{0}_{m},0,\cdots,0)^{\sf T}\quad\mbox{with}\quad\sum_{j=1}^{m}v^{0}_{j}=0.$
Such vectors make $m-1$ dimensional space of eigenvectors giving the zero-eigenvalue $\lambda(m)=0$. Now we take one of such eigenvectors, $\bm{v}^{0}=\bm{v}^{0}_{q}$, $q=1,2,\dots,m-1$, and set
$\bm{v}_{q,\ell}=(S^{\sf T})^{m\ell}\bm{v}^{0}_{q},$
where integers $\ell\in\{0,1,\dots,p_{1}\}$ are chosen so that $\langle\bm{v}_{q,\ell},{\bf 1}\rangle=0$. For example, when $n=6$ and $m=3$, we have $m-1=2$ vectors
$\bm{v}^{0}_{1}=(1,-1,0,0,0,0)^{\sf T},\quad\bm{v}^{0}_{2}=(1,0,-1,0,0,0)^{\sf T}.$
Then, with $p_{1}=[(6-1)/3]=1$ and $\ell\in\{0,1\}$, we obtain the four linearly independent vectors,
$\bm{v}_{1,0}:=\bm{v}^{0}_{1},\quad\bm{v}_{1,1}:=(S^{\sf T})^{3}\bm{v}^{0}_{1}=(0,0,0,1,-1,0)^{\sf T},$
$\bm{v}_{2,0}:=\bm{v}^{0}_{2},\quad\bm{v}_{2,1}:=(S^{\sf T})^{3}\bm{v}^{0}_{2}=(0,0,0,1,0,-1)^{\sf T},$
which satisfy the orthogonality to ${\bf 1}$ and span the generalized eigenspace for the zero-eigenvalue $\lambda(3)=0$ with dimensions $n-p_{1}-1=6-1-1=4$. When $n=5$ and $m=3$, on the other hand, we have also $m-1=2$ vectors
$\bm{v}^{0}_{1}=(1,-1,0,0,0)^{\sf T},\quad\bm{v}^{0}_{2}=(1,0,-1,0,0)^{\sf T}.$
In this case,
$\bm{v}_{2,1}:=(S^{\sf T})^{3}\bm{v}^{0}_{2}=(0,0,0,1,0)^{\sf T}$
does not satisfy the orthogonality condition; $\langle\bm{v}_{2,1},{\bf 1}\rangle\not=0$. Then we have the $n-p_{1}-1=5-[(5-1)/3]-1=5-1-1=3$ dimensional generalized eigenspace spanned by $\{\bm{v}_{1,0},\bm{v}_{1,1},\bm{v}_{2,0}\}$ for the zero-eigenvalue $\lambda(3)=0$. For more systematic study of the generalized eigenspaces of the present models, see [32].
Let $\mathbb{T}_{r}$ and $\mathbb{D}_{r}$ be the circle (one-dimensional torus) and the open disk centered at the origin with radius $r>0$, respectively; $\mathbb{T}_{r}:=\{z\in\mathbb{C};|z|=r\}$ and $\mathbb{D}_{r}:=\{z\in\mathbb{C};|z|<r\}$. We can prove the following by Rouché’s theorem (see, for instance, [1, Section 4.4]).
Proposition 3.3
(i)   All $p_{1}+1$ non-zero eigenvalues, which are given by the solutions of (3.5), lie inside $\mathbb{T}_{n\delta+1}$.
(ii)   Assume that $n\delta>3+2\sqrt{2}=5.82\cdots$. Then the quadratic equation, $r^{2}-(n\delta+1)r+2n\delta=0$, has two real solutions,
$r_{\pm}:=\frac{n\delta}{2}+\frac{1}{2}\pm\frac{n\delta}{2}\sqrt{1-\frac{6}{n\delta}+\frac{1}{(n\delta)^{2}}},$
where $r_{+}=n\delta-1-2/(n\delta)+{\rm O}((n\delta)^{-2})$ and $r_{-}=2+2/(n\delta)+{\rm O}((n\delta)^{-2})$ as $n\delta\to\infty$. The following holds.
(iia)   Only one eigenvalue exists in $\mathbb{D}_{n\delta+1}\setminus\mathbb{D}_{r_{+}}=\{z\in\mathbb{C};r_{+}\leq|z|<n\delta+1\}$.
(iib)   There is no eigenvalue in $\mathbb{D}_{r_{+}}\setminus(\mathbb{D}_{r_{-}}\cup\mathbb{T}_{r_{-}})=\{z\in\mathbb{C};r_{-}<|z|<r_{+}\}$.
(iic)   Other $p_{1}$ non-zero eigenvalues lie in $\mathbb{D}_{r_{-}}\cup\mathbb{T}_{r_{-}}=\{z\in\mathbb{C};|z|\leq r_{-}\}$.
Proof
(i)   We set
$f(z)=z^{p_{1}+1},\quad g(z)=-n\delta\sum_{k=0}^{p_{1}}\left\{1-(p_{1}-k)\frac{m}{n}\right\}z^{k}.$
For $k\in\{0,1,\dots,p_{1}\}$, we see that
$\frac{1}{n}\leq\left|1-(p_{1}-k)\frac{m}{n}\right|\leq 1.$
On $\mathbb{T}_{n\delta+1}$, $|f(z)|=(n\delta+1)^{p_{1}+1}$ and
$|g(z)|\leq n\delta\sum_{k=0}^{p_{1}}|z|^{k}=n\delta\sum_{k=0}^{p_{1}}(n\delta+1)^{k}=(n\delta+1)^{p_{1}+1}-1<|f(z)|.$
By Rouché’s theorem, the numbers of zeros of $f(z)$ and $f(z)+g(z)$ inside $\mathbb{T}_{n\delta+1}$ are the same. Therefore, the assertion is proved.
(ii)   Next we set
$f(z)=-n\delta z^{p_{1}},\quad g(z)=z^{p_{1}+1}-n\delta\sum_{k=0}^{p_{1}-1}\left\{1-(p_{1}-k)\frac{m}{n}\right\}z^{k}.$
Assume $r>1$. Then on $\mathbb{T}_{r}$ we see that
$|g(z)|\leq r^{p_{1}+1}+n\delta\sum_{k=0}^{p_{1}-1}r^{k}=n\delta\left(\frac{r}{n\delta}+\frac{1}{r-1}\right)r^{p_{1}}-\frac{n\delta}{r-1}.$
If
$\frac{r}{n\delta}+\frac{1}{r-1}\leq 1\,\iff\,r^{2}-(n\delta+1)r+2n\delta\leq 0,$ (3.11)
then $|g(z)|<|f(z)|$, and thus the number of solution of (3.5) inside $\mathbb{T}_{r}$ is $p_{1}$. If $n\delta>3+2\sqrt{2}$, then $\mathcal{D}:=(n\delta+1)^{2}-8n\delta>0$ and $r_{\pm}=(n\delta+1\pm\sqrt{\mathcal{D}})/2\in\mathbb{R}$ with $r_{-}<r_{+}$. Hence, if $r_{-}<r<r_{+}$, then (3.11) is satisfied. It is easy to verify that $1<r_{-}<r_{+}<n\delta$. Thus the assertions (iia)–(iic) are proved. ∎
We write the eigenvalue specified by Proposition 3.3 (iia) as $\lambda_{1}(m)$. Let $C_{k},k\in\mathbb{N}_{0}:=\{0,1,2,\dots\}$ be the Catalan numbers, $1,1,2,5,14,42,132,429,\cdots$; that is,
$C_{k}:=\frac{1}{k+1}\binom{2k}{k}=\frac{(2k)!}{(k+1)!k!},\quad k\in\mathbb{N}_{0}.$ (3.12)
See, for instance, [35].
Proposition 3.4
The following holds for $\lambda_{1}(m)$, $1\leq m\leq n$.
(i)   If $p_{1}=0\iff m=n$, then
$\lambda_{1}(n)=n\delta.$ (3.13)
(ii)   If $p_{1}\geq 1\iff m\leq n-1$, then, for $n\delta>1$, we have the expression
$\lambda_{1}(m)$ $=n\delta+1-\sum_{k=0}^{p_{1}-1}C_{k}\left(\frac{m}{n}\right)^{k+1}\frac{1}{(n\delta)^{k}}+{\rm O}((n\delta)^{-p_{1}}).$ (3.14)
Proof
(i)   When $m=n\iff p_{1}=0$, (3.4) is reduced to $z-n\delta=0$ and (3.13) is concluded.
(ii)   When $p_{1}=1$, (3.5) becomes the quadratic equation
$z^{2}-n\delta z-n\delta\left(1-\frac{m}{n}\right)=0,$ (3.15)
and we find
$\lambda_{1}(m)=\frac{n\delta}{2}\left[1+\sqrt{1+\frac{4}{n\delta}\left(1-\frac{m}{n}\right)}\right]=n\delta+1-\frac{m}{n}+{\rm O}((n\delta)^{-1}).$
When $p_{1}\geq 2$, the assertion is proved by induction: For $q=1,2,\dots,p_{1}-1$, we assume
$z=n\delta+1-\sum_{k=0}^{q-1}C_{k}\left(\frac{m}{n}\right)^{k+1}\frac{1}{(n\delta)^{k}}-c\left(\frac{m}{n}\right)^{q+1}\frac{1}{(n\delta)^{q}}$
with unknown coefficient $c$. Insert this into (3.5). Then $c$ is determined to be the $q$-th Catalan number $C_{q}$. ∎
Remark 3.2
(i)   If we set $n=200$, $\delta=0.01$, and $m=1$, (3.14) gives
$\lambda_{1}(1)=2+1-\frac{1}{200}-\frac{1}{200^{2}\times 2}-\frac{2}{200^{3}\times 2^{2}}-\cdots=2.994\cdots.$
This implies the fact that a dot near $z=3$ in Fig. 2 shows an exact eigenvalue. A unit circle with a gap at $z=1$ in this figure shall consist of $p_{1}=[(200-1)/1]=199$ exact eigenvalues as asserted by Theorem 3.2.
(ii)   When $n=200$, we have $p_{1}=[(n-1)/m]=99,24,13$, and 2 for $m=2,8,15$, and 80, respectively. Hence the dots on the outer circle in Fig. 3 and the dots on the outer circle in Fig. 3 are exact eigenvalues. The 13 and the 2 dots, which are not equal to zero in Fig. 3 and Fig. 3, respectively, represent exact eigenvalues. Notice that in these figures the eigenvalue $\lambda_{1}(m)\simeq 3$ is out of the frames. All other $n-p_{1}-1$ eigenvalues are degenerated at the origin.
(iii)   The dots forming the inner circle with radius $\simeq 0.7$ and the two non-zero dots near the origin in Fig. 3, the annulus with wavy boundaries and the two non-zero dots near the origin in Fig. 3, and the small annulus surrounding the origin in Fig. 3 are all not exact eigenvalues. They represent some structures of pseudospectra.
Remark 3.3   If
$1\leq\frac{n}{m}<2\quad\iff\quad\frac{n}{2}<m\leq n,$
then $p_{1}=1$. Assume that $\delta\in\mathbb{R}$. In this case, (3.5) becomes the quadratic equation (3.15). Put $z=x+iy$, $x,y\in\mathbb{R}$. Then we obtain the following equations from (3.15),
$x^{2}-y^{2}-n\delta\left\{x+\left(1-\frac{m}{n}\right)\right\}=0,$
$y(2x-n\delta)=0.$ (3.16)
The second equation in (3.16) gives $y=0$ or $x=n\delta/2$. If we assume $x=n\delta/2$, then the first equation in (3.16) gives $y^{2}=-(n\delta)^{2}/4-n\delta(1-m/n)$. Since $1-m/n\geq 0$, the RHS is negative, and thus this contradicts $y\in\mathbb{R}$. Hence $y=0$. Then the first equation in (3.16) becomes
$x^{2}-n\delta x-n\delta\left(1-\frac{m}{n}\right)=0,$
which is solved by
$x_{\pm}:=\frac{n\delta}{2}\left\{1\pm\sqrt{1+\frac{4}{n\delta}\left(1-\frac{m}{n}\right)}\right\}.$
When $n\delta\gg 1$, the above solutions behave as
$x_{\pm}\simeq\frac{n\delta}{2}\left[1\pm\left\{1+\frac{2}{n\delta}\left(1-\frac{m}{n}\right)\right\}\right]=\begin{cases}\displaystyle{n\delta+1-\frac{m}{n}},\cr\displaystyle{-1+\frac{m}{n}}.\end{cases}$
Hence $x_{+}$ should be identified with $\lambda_{1}(m)$. We see that
$\lim_{m\searrow n/2}x_{-}=-\frac{n\delta}{2}\left\{\sqrt{1+\frac{2}{n\delta}}-1\right\}=:x_{-}^{0}<0,\quad\lim_{m\nearrow n}x_{-}=0.$
Therefore, the non-zero eigenvalue $x_{-}=x_{-}(m)$, which is different from $\lambda_{1}(m)$, moves from $x_{-}^{0}$ to 0 on $\mathbb{R}_{-}$ as $m$ increases from $n/2$ to $n$.
3.2 Model 2
For model 2, we consider the eigenvalue problem,
$S^{(2)}_{\delta J}(m)\bm{v}(m)=\lambda(m)\bm{v}(m).$
If we consider the zero-eigenvalue $\lambda(m)=0$, this equation becomes $(S^{m}+aS^{m+2})\bm{v}(m)=-\delta\alpha(m){\bf 1}$ by (3.7). Since $S^{m}$ and $S^{m+1}$ shift the elements of any vector upward by $m$ and $m+1$ respectively when $S^{m}+aS^{m+1}$ is operated on the vector from the left, the last $m$ elements of the vector $(S^{m}+aS^{m+1})\bm{v}(m)$ are zero. Since ${\bf 1}$ is the all-ones vector, $\alpha(m)$ should be 0. For non-zero eigenvalues $\lambda(m)\not=0$ on the other hand, we can assume $\alpha(m)\not=0$.
For $k\in\mathbb{N}_{0}$, we see that
$\langle(S^{m}+aS^{m+1})^{k}{\bf 1},{\bf 1}\rangle$ $=\left\langle\sum_{q=0}^{k}\binom{k}{q}S^{(k-q)m}a^{q}S^{q(m+1)}{\bf 1},{\bf 1}\right\rangle$
$=\sum_{q=0}^{k}\binom{k}{q}a^{q}\langle S^{mk+q}{\bf 1},{\bf 1}\rangle$
$=\sum_{q=0}^{k}\binom{k}{q}a^{q}\{n-(mk+q)\}1_{(1\leq mk+q\leq n-1)},$
where Lemma 3.1 was used. Let
$p_{2}=p(n-1,m+1),$ (3.17)
where $p(n,m)$ was defined by (3.2). Then Theorem 3.2 is generalized as follows.
Theorem 3.5
For $m\in\{1,2,\dots,n\}$, there are $p_{1}+1$ non-zero eigenvalues, which solve the following equation,
$\frac{1+a}{n\delta}\left(\frac{z}{1+a}\right)^{p_{1}+1}-\frac{\displaystyle{1-\left(\frac{z}{1+a}\right)^{p_{1}+1}}}{\displaystyle{1-\frac{z}{1+a}}}$
+$\qquad+\left\{\frac{m}{n}+\frac{a}{(1+a)n}\right\}\frac{1}{\displaystyle{1-\frac{z}{1+a}}}\left\{p_{1}+1-\frac{\displaystyle{1-\left(\frac{z}{1+a}\right)^{p_{1}+1}}}{\displaystyle{1-\frac{z}{1+a}}}\right\}$
$\qquad-1_{(p_{1}\geq p_{2}+1,\,p_{1}\geq(n+1)/(m+1))}\frac{1}{(1+a)^{p_{1}}}\sum_{k=0}^{p_{1}-p_{2}-1}z^{k}\sum_{q=n-m(p_{1}-k)+1}^{p_{1}-k}a^{q}\binom{p_{1}-k}{q}$
×$\hskip 199.16928pt\times\frac{1}{n}[q-\{n-m(p_{1}-k)\}]=0.$ (3.18)
This equation is also written as the polynomial equation
$\left(\frac{z}{1+a}\right)^{p_{1}+1}-\frac{n\delta}{1+a}\sum_{k=0}^{p_{1}}\left[1-(p_{1}-k)\left\{\frac{m}{n}+\frac{a}{(1+a)n}\right\}\right]\left(\frac{z}{1+a}\right)^{k}$
$\qquad-1_{(p_{1}\geq p_{2}+1,\,p_{1}\geq(n+1)/(m+1))}\frac{n\delta}{(1+a)^{p_{1}+1}}\sum_{k=0}^{p_{1}-p_{2}-1}z^{k}\sum_{q=n-m(p_{1}-k)+1}^{p_{1}-k}a^{q}\binom{p_{1}-k}{q}$
×$\hskip 199.16928pt\times\frac{1}{n}[q-\{n-m(p_{1}-k)\}]=0.$ (3.19)
The corresponding eigenvectors satisfy $\alpha(m)\not=0$. All other $n-p_{1}-1$ eigenvalues degenerate at zero. In this case, the corresponding eigenvectors satisfy $\alpha(m)=0$, that is, they are orthogonal to ${\bf 1}$.
Proof   Following the similar calculation in the proof of Theorem 3.2, (3.8) is replaced by
$\alpha(m)=\delta\alpha(m)\sum_{k=0}^{p_{1}}z^{-(k+1)}\sum_{q=0}^{k}\binom{k}{q}a^{q}\{n-(mk+q)\}1_{(1\leq mk+q\leq n-1)}.$
For non-zero eigenvalues, we can assume $\alpha(m)\not=0$. Then we have
$\frac{1}{n\delta}-\sum_{k=0}^{p_{1}}z^{-(k+1)}\sum_{q=0}^{k}\binom{k}{q}a^{q}\left\{1-\left(\frac{m}{n}k+\frac{q}{n}\right)\right\}1_{(1\leq mk+q\leq n-1)}=0.$
The condition $mk+q\leq n-1$ gives $q\leq n-1-mk$. We see that for $n\in\mathbb{N}$,
$k\leq n-1-mk,\quad m\in\mathbb{N},\quad k\in\mathbb{N}_{0}\quad\iff\quad k\leq p_{2},\quad k\in\mathbb{N}_{0}$
where $p_{2}$ is defined by (3.17). That is, if $k\geq p_{2}+1$, then the summation over $q$ is taken only in the interval $q\in[0,n-1-mk]$ instead of $[0,mk]$. In other words, in the index space $(k,q)\in\mathbb{N}_{0}\times\mathbb{N}_{0}$, the region for which we should take summation is
$\{(k,q);1\leq k\leq p_{2},0\leq q\leq k\}\cup\{(k,q);p_{2}+1\leq k\leq p_{1},0\leq q\leq n-1-mk\}.$
This is equivalent with
$\{(k,q);1\leq k\leq p_{1},1\leq q\leq k\}\setminus\{(k,q);p_{2}+1\leq k\leq p_{1},n-mk\leq q\leq k\}.$
Hence the above equation is written as
$\frac{1}{n\delta}-\sum_{k=0}^{p_{1}}z^{-(k+1)}\sum_{q=0}^{k}\binom{k}{q}a^{q}\left\{1-\left(\frac{m}{n}k+\frac{q}{n}\right)\right\}$
$\quad+\sum_{k=p_{2}+1}^{p_{1}}z^{-(k+1)}\sum_{q=n-mk}^{k}\binom{k}{q}a^{q}\left\{1-\left(\frac{m}{n}k+\frac{q}{n}\right)\right\}=0.$ (3.20)
We notice that
$I_{1}(k)$ $:=\sum_{q=0}^{k}\binom{k}{q}a^{q}\left\{1-\left(\frac{m}{n}k+\frac{q}{n}\right)\right\}$
$=\left(1-\frac{m}{n}k\right)\sum_{q=0}^{k}\binom{k}{q}a^{q}-\frac{1}{n}\sum_{q=0}^{k}q\binom{k}{q}a^{q}$
$=\left(1-\frac{m}{n}k\right)(1+a)^{k}-\frac{1}{n}ak(1+a)^{k-1}$
$=(1+a)^{k}-\frac{1}{n}\frac{(1+a)m+a}{1+a}k(1+a)^{k},$
and hence
$J_{1}$ $:=\sum_{k=0}^{p_{1}}z^{-(k+1)}I_{1}(k)$
$=(1+a)^{-1}\left(\frac{z}{1+a}\right)^{-1}\sum_{k=0}^{p_{1}}\left(\frac{z}{1+a}\right)^{-k}$
$\quad-\frac{1}{n}\frac{(1+a)m+a}{1+a}(1+a)^{-1}\left(\frac{z}{1+a}\right)^{-1}\sum_{k=0}^{p_{1}}k\left(\frac{z}{1+a}\right)^{-k}.$
We apply the summation formulas (3.10) and obtain the expression
$J_{1}$ $=\frac{1}{1+a}\left(\frac{z}{1+a}\right)^{-(p_{1}+1)}\frac{\displaystyle{1-\left(\frac{z}{1+a}\right)^{p_{1}+1}}}{\displaystyle{1-\frac{z}{1+a}}}$
$\qquad-\frac{1}{n}\frac{(1+a)m+a}{(1+a)^{2}}\frac{\displaystyle{\left(\frac{z}{1+a}\right)^{-(p_{1}+1)}}}{\displaystyle{1-\frac{z}{1+a}}}\left\{p_{1}+1-\frac{\displaystyle{1-\left(\frac{z}{1+a}\right)^{p_{1}+1}}}{\displaystyle{1-\frac{z}{1+a}}}\right\}.$
We also see that
$J_{2}:=\sum_{k=p_{2}+1}^{p_{1}}z^{-(k+1)}\sum_{q=n-mk}^{k}a^{q}\binom{k}{q}a^{q}\left\{1-\left(\frac{m}{n}k+\frac{q}{n}\right)\right\}$
$=-\left(\frac{z}{1+a}\right)^{-(p_{1}+1)}\frac{1}{(1+a)^{p_{1}+1}}\sum_{k=p_{2}+1}^{p_{1}}z^{p_{1}-k}\sum_{q=n-mk+1}^{k}a^{q}\binom{k}{q}\frac{1}{n}\{q-(n-mk)\}$
$=-\left(\frac{z}{1+a}\right)^{-(p_{1}+1)}\frac{1}{(1+a)^{p_{1}+1}}\sum_{\ell=0}^{p_{1}-p_{2}-1}z^{\ell}\sum_{q=n-m(p_{1}-\ell)+1}^{p_{1}-\ell}a^{q}\binom{p_{1}-\ell}{q}\frac{1}{n}[q-\{n-m(p_{1}-\ell)\}],$
where we have noticed that the term with $q=n-mk$ vanishes in the first line, and then the summation over $k$ is replaced by the summation over $\ell:=p_{1}-k$ in the last line. $J_{2}$ should be zero if $p_{1}=p_{2}$. For $10\leq\ell\leq p_{1}-p_{2}-$ 1, $n-mp_{1}+1\leq n-m(p_{1}-\ell)+1\leq q\leq p_{1}-\ell\leq p_{1}$ in the second summation in the last line. Hence, if $n-mp_{1}+1>p_{1}\iff p_{1}<(n+1)/(m+1)$, then $J_{2}$ should be zero. Equation (3.20) is thus given by
$\frac{1}{n\delta}-J_{1}+1_{(p_{1}\geq p_{2}+1,\,p_{1}\geq(n+1)/(m+1))}J_{2}=0.$
Multiply the factor $(1+a)\times(z/(1+a))^{p_{1}+1}$. Then (3.18) is obtained. It is easy to verify that this equation is also written as (3.19). ∎
Remark 3.4   The equations (3.18) and (3.19) are also written as follows,
$\frac{z^{p_{3}+1}}{n\delta}-\left\{1-\frac{z}{z-(1+a)}\left(m+\frac{1}{2}\right)\frac{1}{n}\right\}\frac{z^{p_{3}+1}-(1+a)^{p_{3}+1}}{z-(1+a)}$
$\qquad-\left(m+\frac{1}{2}\right)\frac{1}{n}\frac{z^{p_{3}+1}+p_{3}(1+a)^{p_{3}+1}}{z-(1+a)}$
$\qquad-1_{(p_{1}\geq p_{3}+1,\,p_{3}\leq(n-m-1)/m)}\sum_{k=p_{3}+1}^{p_{1}}z^{p_{3}-k}\sum_{r=0}^{n-mk-1}\binom{k}{r}\left(1-\frac{m}{n}k-\frac{r}{n}\right)a^{r}=0$ (3.21)
with $p_{3}:=p(n,m+1)$, where $p(n,m)$ was defined by (3.2). As a matter of course, if we put $a=0$, (3.18), (3.19), and (3.21) are reduced to (3.4) and (3.5).
(a) $m=1$
(b) $m=2$
(c) $m=3$
(d) $m=4$
Figure 5: Exact eigenvalues are plotted for model 2, $(S^{(2)}_{\delta J}(m))_{m=1}^{n}$, with $n=200$, $\delta=0.01$, and $a=1$ at $m=1,2,3$, and 4, respectively.
Remark 3.5 Consider the difference in $m$ of $\widehat{p}$ defined by (3.2),
$\Delta\widehat{p}$ $:=\widehat{p}(n-1,m)-\widehat{p}(n-1,m+1)$
$=\frac{n-1}{m}-\frac{n-1}{m+1}=\frac{n-1}{m(m+1)}.$
Here we regard $\Delta\widehat{p}$ as a function of the real variable $m\geq 1$, and for $s\not=0$ we solve the equation
$\Delta\widehat{p}=s\quad$ $\iff\quad\frac{n-1}{m(m+1)}=s$
$\iff\quad m^{2}+m-\frac{n-1}{s}=0.$
Let $m(n,s)$ be the positive solution,
$m(n,s):=\sqrt{\frac{n-1}{s}+\frac{1}{4}}-\frac{1}{2}.$ (3.22)
By definition, if $1\leq m<m(n,s)$, then $\Delta\widehat{p}>s$ $\Longrightarrow p_{1}-p_{2}\gtrsim s$, and if $m(n,1/s)<m$, then $\Delta\widehat{p}<1/s$ $\Longrightarrow p_{1}-p_{2}\lesssim 1/s$. Suppose that $s\in\mathbb{N}$. Then the above calculation will be interpreted as follows: When $m\simeq m(n,s)$, the number of terms in the last part of the left-hand-side of (3.18) is about $s$. And at about $s$ successive values of $m$ around $m(n,1/s)$, $p_{1}=p_{2}$, and hence the last part of the left-hand-side of (3.18) vanishes. For example, when $n=10^{5}$, $m(10^{5},10)=99.50\cdots$, $m(10^{5},1/10)=999.4\cdots$, and we can see that $p_{1}-p_{2}=999-990=9$ at $m=100$, and that $p_{1}=p_{2}=99$ for the ten values of $m$; $m=1000,1001,\cdots,1009$.
The dependence of $m(m,s)$ on $n$ expressed by $\sqrt{n-1}$ in (3.22) is essential, and the following lemma is valid. We write the smallest integer greater than or equal to $x\in\mathbb{R}$ as $\lceil x\rceil$ (the ceiling function of $x$). Remark that the floor function of $x$ is denoted by $[x]$ in this paper.
Lemma 3.6
Let $I_{n-1}:=[\lceil\sqrt{n-1}\,\rceil,n-1]\cap\mathbb{N}$ and $T_{n-1}:=\{[(n-1)/k];k=1,2,\dots,n-1\}$. Then
$p_{1}-p_{2}=\begin{cases}1,\quad\mbox{if $m\in I_{n-1}\cap T_{n-1}$},\cr 0,\quad\mbox{if $m\in I_{n-1}\setminus T_{n-1}$}.\end{cases}$
The proof is given in Appendix A. Notice that $\lceil\sqrt{n-1}\,\rceil$ in the definition of $I_{n-1}$ can not be replaced by $[\sqrt{n-1}\,]$ for general $n$. For instance, when $n=9$, $[\sqrt{9-1}\,]=[2.8\cdots]=2$, while $\lceil\sqrt{9-1}\,\rceil=3$, and $p_{1}-p_{2}=[8/2]-[8/3]=2$ for $m=2$. When $n=200$, $\lceil\sqrt{n-1}\,\rceil=\lceil 14.1\cdots\rceil=15$. Since $[(n-1)/13]=[15.3\cdots]=15$, $I_{n-1}\cap T_{n-1}=\{[199/k];k=1,2,\dots,13\}$.
Proposition 3.4 for model 1 is generalized for model 2 as follows.
Proposition 3.7
We have an eigenvalue $\lambda_{1}(m)$, whose modulus goes to $\infty$ as $n\delta\to\infty$. The following holds for $m\in\{1,2,\dots,n\}$.
(i)   If $p_{1}=0\iff m=n$, then
$\lambda_{1}(n)=n\delta.$
(ii)   If $p_{1}\geq 1\iff m\leq n-1$, then, for $n\delta>1$, we have the expression
$\lambda_{1}(m)$ $=n\delta+1+a-(1+a)\sum_{k=0}^{p_{1}-1}C_{k}\left(\frac{m}{n}+\frac{a}{(1+a)n}\right)^{k+1}\left(\frac{1+a}{n\delta}\right)^{k}+{\rm O}((n\delta)^{-p_{1}})$
$\qquad\qquad+1_{(p_{1}\geq p_{2}+1)}{\rm O}((n\delta)^{-p_{2}}).$ (3.23)
Proof   Comparing the left-hand-side of (3.4) for model 1 and the first two parts in the left-hand-side of (3.18) for model 2, we find that the latter is obtained from the former by the following replacement,
$z\to\frac{z}{1+a},\quad n\delta\to\frac{n\delta}{1+a},\quad\frac{m}{n}\to\frac{m}{n}+\frac{a}{(1+a)n},$ (3.24)
associated with the introduction of the parameter $a$ in model 2. The first line of (3.23) is obtained from (3.14) by this replacement (3.24). The correction ${\rm O}((n\delta)^{-p_{2}})$ should be added due to the last part in the left-hand-side of (3.18). ∎
Remark 3.6
(i)   If we set $n=200$, $\delta=0.01$, and $a=1$, (3.23) gives
$\lambda_{1}(m)\simeq 2+1+1-2\left(\frac{m}{200}+\frac{1}{2\times 200}\right)=3.995-\frac{m}{100}.$
This implies the fact that a dot near $z=4$ in each figure of Fig. 4 shows an exact eigenvalue.
(ii)   The exact eigenvalues given by the solutions of (3.18) of Theorem 3.5 are plotted in Fig. 5 for $m=1,2,3$, and 4. Comparing Fig. 4 and Fig. 5, the dots located in the outermost regions are exact eigenvalues. The exact eigenvalues located in the inner regions, especially most of the exact eigenvalues on $\mathbb{R}_{-}$, are missing in the numerical results. The patterns observed in the vicinity of the origin in the numerical results (Fig. 4) represent the structures of pseudospectra.
4 Pseudospectrum Processes
Consider the banded Toeplitz matrices such that the number of diagonal lines in which the elements are non-zero is finite and given by $12w+$ 1, $w\in\mathbb{N}_{0}$. Let $\{A_{n}\}$ be a family of such Toeplitz matrices with sizes $n\in\mathbb{N}$;
$A_{n}=\Big{(}(A_{n})_{jk}\Big{)}_{1\leq j,k\leq n}=\Big{(}a_{j-k}1_{(|j-k|\leq w)}\Big{)}_{1\leq j,k\leq n}.$
We write the matrix representation of the corresponding banded Toeplitz operator as
$\widehat{A}=\Big{(}a_{j-k}1_{(|j-k|\leq w)}\Big{)}_{j,k\in\mathbb{N}}.$
The symbol of $\widehat{A}$ is defined as [11, 39]
$f_{\widehat{A}}(z):=\sum_{\ell;|\ell|\leq w}a_{\ell}z^{\ell}.$
Let $\mathbb{T}:=\mathbb{T}_{1}=\{e^{i\theta};\theta\in[0,2\pi)\}$, i.e. the unit circle (the one-dimensional unit torus). Then the symbol curve [11, 39] is defined by
$f_{\widehat{A}}(\mathbb{T})=\{f_{\widehat{A}}(z);z\in\mathbb{T}\}.$
Given a point $z\in\mathbb{C}\setminus f_{\widehat{A}}(\mathbb{T})$, the winding number $I(f_{\widehat{A}},z)$ is defined as the winding number of $f_{\widehat{A}}(\mathbb{T})$ about $z$ in the usual positive (counterclockwise) sense. The following theorem is well-known [11, 39].
Theorem 4.1
Let $\sigma(\widehat{A})$ be the spectra of $\widehat{A}$. Then $\sigma(\widehat{A})$ is equal to $f_{\widehat{A}}(\mathbb{T})$ together with all the points enclosed by this curve with $I(f_{\widehat{A}},z)\not=0$.
The following fact was proved [34, 39].
Proposition 4.2
For some $M>1$ and all sufficiently large $n$,
$\|(zI-A_{n})^{-1}\|\geq M^{n}\quad\mbox{for any $z\in\sigma(\widehat{A})$}.$
This implies that the pseudospectra of $A_{n}$ will reflect the exact spectra of the corresponding Toeplitz operator $\widehat{A}$.
4.1 Model 1
(a) $m=1$
(b) $m=2$
(c) $m=3$
(d) $m=4$
(e) Color scale
Figure 6: Contour plots of $\|(zI-S^{(1)}_{\delta J}(m))^{-1}\|$ for model 1 with $n=50$ and $\delta=0.01$ at $m=1,2,3$, and 4, respectively.
The Toeplitz operators corresponding to $S^{m}$, $m\in\{1,2,\dots,n\}$, are given by
$\widehat{S}^{m}=\Big{(}(\widehat{S}^{m})_{jk}\Big{)}_{j,k\in\mathbb{N}}=\Big{(}\delta_{j\,k-m}\Big{)}_{j,k\in\mathbb{N}},\quad m\in\mathbb{N}.$
The symbols of $\widehat{S}^{m}$, $m\in\mathbb{N}$ are given by
$f_{\widehat{S}^{m}}(z)=z^{m}.$
Let $\mathbb{D}$ be the open unit disk, $\mathbb{D}:=\mathbb{D}_{1}=\{z\in\mathbb{C};|z|<1\}$, and we write $\overline{\mathbb{D}}:=\{z\in\mathbb{C};|z|\leq 1\}=\mathbb{D}\cup\mathbb{T}$. We see that $f_{\widehat{S}^{m}}(\mathbb{T})=\mathbb{T}$ and all points enclosed by $\mathbb{T}$ have winding number $m\in\mathbb{N}\not=0$. Hence by Theorem 4.1,
$\sigma(\widehat{S}^{m})=\overline{\mathbb{D}},\quad m\in\mathbb{N}.$
The present model 1 can be regarded as the system such that the rank 1 perturbation $\delta J$ is added to $S^{m}$, $m=1,2,\dots,n$. It has been reported in many examples (see, for instance, Fig. 7.4 and explanations in Section 7 of [39]), caused by dense random perturbations the eigenvalues of nonnormal Toeplitz matrices tend to ‘trace out’ strikingly the pseudospectra of non-perturbed matrices. As suggested by Proposition 4.2, the boundaries of the pseudospectra will be lined up along the symbol curves of the corresponding Toeplitz operators. Mathematical study of random perturbations of banded Toeplitz matrices, see [5, 6, 7, 8, 9, 36] and references therein.
We notice that since $J$ in the present perturbation-term $\delta J$ is the all-ones matrix, the Toeplitz operators corresponding to $S^{(1)}_{\delta J}(m)$, $m=1,2,\dots,n$, are not banded and hence symbols can not be defined. We consider, however, that the inner circle and annuli found in the numerical results, Figs. 33, represent the boundaries and some structures of the $\varepsilon$-pseudospectra with appropriate values of $\varepsilon$.
Figure 6 shows the contour plots of the 2-norms of the resolvents, $\|(zI-S^{(1)}_{\delta J}(m))^{-1}\|$, for $m=1,2,3$, and 4, where $n=50$ and $\delta=0.01$. Here the dots in the outer regions denote the exact eigenvalues. The values of $\|(zI-S^{(1)}_{\delta J}(m))^{-1}\|$ grow exponentially up to $10^{5}$ as we approach to the origin. The $\varepsilon$-pseudospectrum with a given small value of $\varepsilon$ decreases monotonically as $m$ increases.
4.2 Model 2
(a) $m=1$
(b) $m=2$
(c) $m=3$
(d) $m=4$
Figure 7: The symbol curves of $S^{m}+aS^{m+1}$ with $a=1$ at $m=1,2,3$, and 4.
(a) $m=1$
(b) $m=2$
(c) $m=3$
(d) $m=4$
Figure 8: The inner parts of the numerically obtained eigenvalues for model 2 showing the structures of pseudospectra.
Let $\widehat{S}^{m}+a\widehat{S}^{m+1}$, $m\in\mathbb{N}$ be the Toeplitz operators corresponding to $S^{m}+aS^{m+1}$, $m\in\{1,2,\dots,n\}$. The symbols are given by
$f_{\widehat{S}^{m}+a\widehat{S}^{m+1}}(z)=z^{m}+az^{m+1},\quad m\in\mathbb{N}.$
In Fig. 8, we show the symbol curves $f_{\widehat{S}^{m}+a\widehat{S}^{m+1}}(\mathbb{T})$ with $a=1$ for $m=1,2,3$, and 4. We can consider that each symbol curve consists of $m+1$ closed simple curves with different sizes. Each closed simple curve is symmetric with respect to the real axis $\mathbb{R}$ and adjacent ones osculate at a point on $\mathbb{R}$. For instance, in the symbol curve for $m=3$ in Fig. 8, the osculating points of the composing closed simple curves are at about $-1.8,1.2$, and $-0.4$. We consider to separate them into two parts; a single outermost closed simple curve and an inner part composed of $m$ smaller closed simple curves osculating each other.
Now we consider such separation also for the plots of numerically obtained eigenvalues shown in Fig. 4. We have observed the following.
(i)   At each time $m$, the dots composing the outermost curve shown in the numerical result, Fig. 4, coincide with the dots in the outermost curve consisting of the exact eigenvalues shown in Fig. 5.
(ii)   We cut out the inner parts composed by dots in the numerical result, Fig. 4, and show them in Fig. 8. They seem to be the scale-downs of the inner parts of the symbol curves $f_{\widehat{S}^{m}+a\widehat{S}^{m+1}}(\mathbb{T})$ shown in Fig. 8. Such structures can not be found in Fig. 5 plotting the exact eigenvalues. The size of the inner part decreases rapidly as $m$ increases and the complicated patterns of the inner parts are smeared out for large values of $m$.
(iii)   Many of the exact eigenvalues on $\mathbb{R}_{-}$ found in Fig. 5 are hidden inside of the inner parts. Additional dots, which are not found in Fig. 5, are observed only inside the inner parts.
Based on the above observations, we give the following conjecture for model 2.
Conjecture 4.3
For model 2, the following holds.
(i)   At each time $m$, the dots composing the outermost curve shown in the numerical result, Fig. 4, are exact eigenvalues of $S_{\delta J}^{(2)}(m)$.
(ii)   The inner parts composed by dots in the numerical result, Fig. 4, are not exact eigenvalues of $S_{\delta J}^{(2)}(m)$ for $m\geq 2$. They are the eigenvalues of the perturbed system of $S_{\delta J}^{(2)}(m)$ in which uncontrolled rounding errors of computer are added. The distributions of the eigenvalues of such perturbed system represent the structures of pseudospectrum including the zero-eigenvalue of the original system, which reflect the inner parts of the spectra of the Toeplitz operator, $\sigma(\widehat{S}^{m}+a\widehat{S}^{m-1})$, without the deterministic perturbation by $\delta J$ nor uncontrolled perturbations by rounding errors of computer. The size of the pseudospectrum including the zero-eigenvalue, which is represented by the inner part of symbol curve, decreases exponentially as $m$ increases.
5 Asymptotics in Infinite-Matrix Limits
For model 1, we fix $m$ and $\delta$ so that they satisfy $m\leq n-1$ and $\delta>4m/n^{2}$. Proposition 3.4 (ii) implies
$\lambda_{1}(m)\sim\frac{n\delta}{2}+1+\frac{n\delta}{2}\sqrt{1-\frac{4m}{n^{2}\delta}}=n\delta+1-\frac{m}{n}+{\rm O}(n^{-2})\to\infty\quad\mbox{as $n\to\infty$}.$ (5.1)
Here we have used the formula of the generating function for the Catalan numbers [35],
$\sum_{k=0}^{\infty}C_{k}x^{k}=\frac{1-\sqrt{1-4x}}{2x}\quad\mbox{for}\,\,x<\frac{1}{4}.$
The fact (5.1) implies that $\lambda_{1}(m)$ solves the quadratic equation asymptotically,
$z^{2}-(n\delta+2)z+1+n\delta\left(1+\frac{m}{n}\right)=0\quad\mbox{as $n\to\infty$}.$ (5.2)
We write the solution of (5.2) other than (5.1) as $\widetilde{\lambda}_{1}(m)$, which behaves as
$\widetilde{\lambda}_{1}(m)=\frac{n\delta}{2}+1-\frac{n\delta}{2}\sqrt{1-\frac{4m}{n^{2}\delta}}=1+\frac{m}{n}+{\rm O}(n^{-2})\to 1\quad\mbox{as $n\to\infty$}.$ (5.3)
Then it is easy to verify that the left-hand-side of Eq. (3.5) in Theorem 3.2 is written as
$z^{p_{1}+1}-n\delta\sum_{k=0}^{p_{1}}\left\{1-(p_{1}-k)\frac{m}{n}\right\}z^{k}$
$=(z-\lambda_{1}(m))(z-\widetilde{\lambda}_{1}(m))\sum_{k=0}^{p_{1}-1}(p_{1}-k)z^{k}+(p_{1}+1)\left\{z-\left(n\delta+1-\frac{1}{p_{1}+1}\right)\right\}.$ (5.4)
We notice that
$\frac{p_{1}+1}{p_{1}}\left\{z-\left(n\delta+1-\frac{1}{p_{1}+1}\right)\right\}=z-(n\delta+1)-\frac{n\delta}{p_{1}}+{\rm O}(n^{-1})$
=$\qquad=\left\{z-\left(n\delta+1-\frac{m}{n}+{\rm O}(n^{-2})\right)\right\}\left(1+\frac{1}{p_{1}}+{\rm O}(n^{-2})\right)$
$\qquad\sim(z-\lambda_{1}(m))\left(1+\frac{1}{p_{1}}\right)\quad\mbox{as $n\to\infty\,\,(p_{1}\to\infty)$}$
for (5.1). Then (5.4) is factorized by $z-\lambda_{1}(m)$ asymptotically in the sense that
$\frac{1}{p_{1}}\left[z^{p_{1}+1}-n\delta\sum_{k=0}^{p_{1}}\left\{1-(p_{1}-k)\frac{m}{n}\right\}z^{k}\right]$
$\,\sim(z-\lambda_{1}(m))\left\{(z-\widetilde{\lambda}_{1}(m))\sum_{k=0}^{p_{1}-1}\left(1-\frac{k}{p_{1}}\right)z^{k}+1+\frac{1}{p_{1}}\right\}\quad\mbox{as $n\to\infty\,\,(p_{1}\to\infty)$}.$ (5.5)
Proposition 5.1
For model 1, fix $m$ and $\delta$ satisfying $m\leq n-1$ and $\delta>4m/n^{2}$. Then as $n\to\infty$, $p_{1}$ non-zero exact eigenvalues except $\lambda_{1}(m)$ tend to be well approximated by
$e^{2\pi i\ell/(p_{1}+1)},\quad\ell=1,2,\dots,p_{1}.$
That is, the eigenvalues become to form a configuration such that one point at $z=1$ is eliminated from the equidistance $p_{1}+1$ points $\{e^{2\pi i\ell/(p_{1}+1)};\ell=0,1,\dots,p_{1}\}$ on $\mathbb{T}$.
Proof   For (5.3) $\widetilde{\lambda}_{1}(m)\in\mathbb{R}$ and $\widetilde{\lambda}_{1}(m)\to 1$ as $n\to\infty$. Eq. (5.5) including the term $1+1/p_{1}$ in the right-hand-side implies that $\widetilde{\lambda}_{1}(m)$ does not satisfy Eq. (3.5) in Theorem 3.2. By (5.3) and the summation formulas (3.10), we see that
$(z-\widetilde{\lambda}_{1}(m))\sum_{k=0}^{p_{1}-1}\left(1-\frac{k}{p_{1}}\right)z^{k}+1+\frac{1}{p_{1}}$
$\quad\sim\left(z-1-\frac{m}{n}\right)\left\{\frac{1}{1-z}-\frac{z(1-z^{p_{1}})}{p_{1}(1-z)^{2}}\right\}+1+\frac{1}{p_{1}}$
$\quad\sim-\frac{m}{n}\frac{1}{1-z}+\frac{z(1-z^{p_{1}})}{p_{1}(1-z)}+\frac{1}{p_{1}}$
$\quad=-\frac{1}{p_{1}(1-z)}\left[z^{p_{1}+1}-\left(1-\frac{m}{n}p_{1}\right)\right]\quad\mbox{as $n\to\infty\,\,(p_{1}\to\infty)$}.$
We write the solution of the equation
$z^{p_{1}+1}=1-\frac{m}{n}p_{1}$
as $z=re^{i\theta}$, $r>0$, $\theta\in[0,2\pi)$. Then we have
$\log r=\frac{1}{p_{1}+1}\log\left(1-\frac{m}{n}p_{1}\right),$
$(p_{1}+1)\theta=0\quad\mbox{mod $2\pi$}.$
They give $r\to 1$ as $n\to\infty$ ($p_{1}\to\infty$) and $\theta=2\pi\ell/(p_{1}+1)$, $\ell=1,2,\dots,p_{1}$. Here the case $\theta=0$ ($\ell=0$) should not be included, since $\widetilde{\lambda}_{1}(m)$ is not the solution as mentioned above. The assertion is hence proved. ∎
Now we consider the quadratic equation for $z$,
$\left(\frac{z}{1+a}\right)^{2}-\left(\frac{n\delta}{1+a}+2\right)\frac{z}{1+a}+1+\frac{n\delta}{1+a}\left\{1+\frac{m}{n}+\frac{a}{(1+a)n}\right\}=0,$ (5.6)
which is obtained from (5.2) by the replacement (3.24). We write the solutions of (5.6) as
$\kappa_{\pm}:=\frac{n\delta}{2}+1+a\pm\frac{n\delta}{2}\sqrt{1-\frac{4\{(1+a)m+a\}}{n^{2}\delta}}.$
Note that
$\kappa_{+}=n\delta+1+a-(1+a)\sum_{k=0}^{\infty}C_{k}\left(\frac{m}{n}+\frac{a}{(1+a)n}\right)^{k+1}\left(\frac{1+a}{n\delta}\right)^{k},$
where $C_{k},k=0,1,\dots$, are the Catalan numbers (3.12). The first line of (3.23) in Proposition 3.7 (ii) for the eigenvalue $\lambda_{1}(m)$ of model 2 is regarded as a truncation of the infinite series of $\kappa_{+}$. Then the first line of the left-hand-side of (3.19) in Theorem 3.5 is written as follows,
$\left(\frac{z}{1+a}\right)^{p_{1}+1}-\frac{n\delta}{1+a}\sum_{k=0}^{p_{1}}\left[1-(p_{1}-k)\left\{\frac{m}{n}+\frac{a}{(1+a)n}\right\}\right]\left(\frac{z}{1+a}\right)^{k}$
$\quad=\left(\frac{z}{1+a}-\frac{\kappa_{+}}{1+a}\right)\left(\frac{z}{1+a}-\frac{\kappa_{-}}{1+a}\right)\sum_{k=0}^{p_{1}-1}(p_{1}-k)\left(\frac{z}{1+a}\right)^{k}$
$\qquad+(p_{1}+1)\left\{\frac{z}{1+a}-\left(\frac{n\delta}{1+a}+1-\frac{1}{p_{1}+1}\right)\right\}.$
This equality can be regarded as the extension of (5.4) including $a$ obtained by the replacement (3.24).
Lemma 3.6 clarified the condition for $m$ so that $p_{1}=p_{2}$; that is, $m\in I_{n-1}\setminus T_{n-1}$. Since $\min I_{n-1}=\lceil\sqrt{n-1}\,\rceil$, if $m\gtrsim\min I_{n-1}$, then $p_{1}=[(n-1)/m]\to\infty$ as $n\to\infty$. Proposition 5.1 is generalized for model 2 in such a situation.
Proposition 5.2
For model 2, fix $\delta$ and $a$. Consider $m$ satisfying $\delta>4\{(1+a)m+a\}/n^{2}$ and $p_{1}=p_{2}$. Then as $n\to\infty$, $p_{1}$ non-zero exact eigenvalues except $\lambda_{1}(m)$ become to be well approximated by
$(1+a)e^{2\pi i\ell/(p_{1}+1)},\quad\ell=1,2,\dots,p_{1}.$
That is, the eigenvalues become to form a configuration such that one point at $z=1+a$ is eliminated from the equidistance $p_{1}+1$ points $\{(1+a)e^{2\pi i\ell/(p_{1}+1)};\ell=0,1,\dots,p_{1}\}$ on $\mathbb{T}_{1+a}$.
(a) $(n,m)=(1000,4)$ with $m/n=1/250$.
(b) $(n,m)=(1000,20)$ with $m/n=1/50$.
Figure 9: Numerically obtained eigenvalues are plotted for model 2, $(S^{(2)}_{\delta J}(m))_{m=1}^{n}$, with $\delta=0.01$ and $a=1$ for different values of $(n,m)$. They shall be compared with Fig. 4 for $(n,m)=(200,4)$ with $m/n=1/50$.
We have performed numerical calculations of model 1 and model 2 for a variety of the matrix size $n$ with $m\in\{1,2,\dots,n\}$. Then we arrived at the following conjectures.
Conjecture 5.3
For model 1, at each fixed $m\geq 1$, the boundary of the pseudospectrum including the zero-eigenvalue increases its size as $n$ increases. It converges to the unit circle $\mathbb{T}=f_{\widehat{S}^{m}}(\mathbb{T})$ as $n\to\infty$. The inside of $\mathbb{T}$ becomes to be fulfilled by the eigenvalues of perturbed system as $n\to\infty$.
Conjecture 5.4
For model 2, the following holds.
(i)   At each fixed $m\geq 1$, the exact eigenvalues composing outermost curve converge to the outermost closed simple curve of the symbol curve $f_{\widehat{S}^{m}+a\widehat{S}^{m+1}}(\mathbb{T})$ as $n\to\infty$. These eigenvalues are insensitive to random perturbations and rounding errors of computer.
(ii)   At each fixed $m\geq 1$, the distribution of eigenvalues of perturbed system observed in the inner part represents the pseudospectrum including the zero-eigenvalue of $S_{\delta J}^{(2)}(m)$. The size of the pseudospectrum including the zero-eigenvalue increases as $n$ increases. It converges to the inner part of the the symbol curve $f_{\widehat{S}^{m}+a\widehat{S}^{m+1}}(\mathbb{T})$ as $n\to\infty$. Only inside of the inner part of the the symbol curve $f_{\widehat{S}^{m}+a\widehat{S}^{m+1}}(\mathbb{T})$ becomes to be fulfilled by the eigenvalues of perturbed systems as $n\to\infty$.
Remark 3.5 and Lemma 3.6 in Section 3.2 suggest that in the period $1\leq m<\lceil\sqrt{n-1}\,\rceil$, $p_{1}\geq p_{2}+1$ and the last part of the left-hand-side of (3.18) should play an important role to determine the exact eigenvalues of model 2. For such $m$, however, the ratio $m/n$ becomes zero as $n\to\infty$. Hence, when we consider the case in which both of $n$ and $m$ are sufficiently large with a fixed value of the ratio $m/n>0$, the last part of the left-hand-side of (3.18) becomes irrelevant to determine the exact eigenvalues. Lemma 3.6 implies that if $m\in I_{n-1}\cap T_{n-1}$, then $p_{1}-p_{2}=1$ and the last part of the left-hand-side of (3.18) gives a term which does not include $z$ but depends on the value of $m$ and the parameter $a$. It is obvious that for each $m=[(n-1)/k]\in I_{n-1}\cap T_{n-1}$, $k=1,2,\dots$, we have $m/n\to 1/k$ as $n\to\infty$. We also notice that the additional term $a/\{(1+a)n\}$ for the last formula in the replacement (3.24) becomes zero as $n\to\infty$ with fixed $a$. From the numerical calculations and the above considerations, we have the following conjecture.
Conjecture 5.5
For both of model 1 and model 2, the sizes of the spectra and the pseudospectra are determined by the ratio $m/n$, if both of $n$ and $m(\leq n)$ are sufficiently large and the other parameters $\delta$ and $a$ are fixed. Hence in the numerically observed eigenvalues with different values of the pair $(n,m)$ but with the same ratio $n/m$, we will observe the patterns with the components which are in the similar sizes with each other, if $n$ and $m(\leq n)$ are both sufficiently large. In other words, we will have nontrivial scaling limits $n\to\infty$ and $m\to\infty$ with fixed $m/n\in(0,1)$.
Figure 9 demonstrates parts of Conjectures 5.4 (ii) and 5.5 for model 2, where we have fixed $\delta=0.01$ and $a=1$. Compared with Fig. 4 for $(n,m)=(200,4)$, the size of the inner part becomes larger in Fig. 9 associated with the increment of the matrix size $n$ from 200 to 1000 with fixed value of $m=4$ (see Conjectures 5.4 (ii)). Notice that the structures in the inner parts both reflect the inner part of the symbol curve of $S^{m}+aS^{m+1}$ with $m=4$ found in Fig. 8. In Fig. 9 for $(n,m)=(1000,20)$, the structure of the inner part becomes much complicated reflecting the symbol curve of $S^{m}+aS^{m+1}$ with $m=20$, but its size is approximately in the same as that in Fig. 4 for $(n,m)=(200,4)$, since the ratio is the same; $m/n=1/50$ (see Conjecture 5.5).
6 Concluding Remarks and Future Problems
Figure 10: (a) Dependence on the ratio $m/n$ of the mean radius $R(m)$ of eigenvalues for the random process $(S_{\delta Z}(m))_{m=1}^{n}$ for $n=200$ and $\delta=0.01$ evaluated by averaging over $10^{3}$ sample processes. (b) Devil’s staircase-like structure observed in the first derivative $R(m+1)-R(m)$ for the random process $(S_{\delta Z}(m))_{m=1}^{n}$. (c) Deformed version of staircase structure observed in the first derivative $R(m+1)-R(m)$ for model 1. (d) Deformed version of staircase structure observed in the first derivative $R(m+1)-R(m)$ for model 2.
Now we discuss a possibility to study the discrete-time random process $(S_{\delta Z}(m))_{m=1}^{n}$ defined as (1.4) by comparing it with the present models. When we numerically simulate the eigenvalue process of $(S_{\delta Z}(m))_{m=1}^{n}$, we observe that the obtained dots form an annulus which seems to be similar to Fig. 1.
The size of annulus, however, rapidly decreases as the discrete-time $m$ increases and the annulus becomes a shrinking disk. We have calculated the mean radius of annulus or disk at each time $m$, which is defined as the mean of radial coordinates of all dots of numerically obtained eigenvalues. We have performed $10^{3}$ independent simulations and averaged the time-dependence of mean radius. As shown by Fig. 10, the expectation of mean radius, $R(m)$, which is evaluated by averaging over $10^{3}$ sample processes, decreases monotonically in time $m$. Since a cusp is observed at $m/n=1/2$, we have calculated the first derivative, that is, the increment $R(m+1)-R(m)$, $m=1,2,\dots,n$, numerically. The result shows devil’s staircase-like structure as shown in Fig. 10, where the thin vertical lines are given at $m/n=1/k$, $k=2,3,\dots,7$ from the right to the left.
As shown by Fig. 3 and Fig. 6 for model 1 and by Fig. 4 for model 2, we have observed the similar monotonic reduction of sizes of exact-spectra and pseudospectra in time $m$. As asserted by Conjecture 5.5, this phenomenon will be described by the ratio $m/n$, when $n$ and $m(\leq n)$ are both sufficiently large. Figs. 10 and 10 show the $m/n$-dependence of the fist derivatives of the mean radii of all numerically observed eigenvalues for model 1 and model 2, respectively. We see deformed but similar staircase structures. In these figures, the thin vertical lines are given at $m/n$ with times $m=[(n-1)/k]$, $k=1,2,\dots,7$, which are included in $I_{n-1}\cap T_{n-1}$ and at which $p_{1}-p_{2}=1$ as proved by Lemma 3.6 in Section 3.2. For $m\in I_{n-1}\cap T_{n-1}$, when $m\to m+1$ the degree $p_{1}+1$ of the polynomial equation for the exact eigenvalues (3.5) of model 1 (resp. (3.19) of model 2) decreases by one, and at time $m$ the last part of the left-hand-side of (3.19) can give an additional constant term to the polynomial equation for model 2. (For instance, when $n=200$, $p_{1}-p_{2}=1$ if $m=[(n-1)/k]=[199/k]$, $k=1,2,\dots,13$ as claimed just after Lemma 3.6. The last part of the left-hand-side of (3.19) does not give any contribution, however, when $k=1,2,4,5,8$, and 10, since the condition $p_{1}\geq(n+1)/(m+1)$ is not satisfied.)
The similarity of Fig. 10 to Figs. 10 and 10 suggests some connection between the present deterministic processes of non-banded Toeplitz matrices and the banded Toeplitz matrices with random perturbations [5, 6, 7, 8, 9, 36].
We list out other future problems.
(i)   In the present paper, we have reported mainly the special case $\delta>0$ in the numerical results, although the theorems and the propositions proved in Sections 3 and 5 are valid for $\delta\in\mathbb{C}$. It is an interesting future problem to study the present models systematically for general values of $\delta\in\mathbb{C}$.
(ii)   Detained analysis of the equations (3.18) and (3.19) including the last parts is needed to understand the exact-eigenvalue processes in the early stage, $1\leq m<\lceil\sqrt{n-1}\,\rceil$. Precise descriptions are desired for the eigenvalues forming the deformed outer circles and the distributions on $\mathbb{R}_{-}$ shown in Fig. 5 for $m=1,2,3$, and 4.
(iii)   We have distinguished the exact non-zero eigenvalues which are insensitive to perturbations from the numerically observed eigenvalues which are not the eigenvalues of the original model $S_{\delta J}^{(\ell)}(m)$, but are the eigenvalues of perturbed systems. The distributions of the latter eigenvalues visualize the pseudospectrum including the zero-eigenvalue of the original model. Mathematical proofs for Conjectures 4.3, 5.3, 5.4, and 5.5 will be challenging future problems. See [32] for further considerations.
We hope that the study of the stochastic processes of eigenvalues and pseudospectra of $(S_{\delta Z}(m))_{m=1}^{n}$ will be useful to understand the phenomena observed by Burda et al. [15] in the non-Hermitian matrix-valued BM, $(M(t))_{t\geq 0}$, started from nonnormal matrices with defective eigenvalues. Notice that the pseudospectra can be evaluated using the condition numbers of eigenvalues (the Bauer–Fike theorem [39, Sects. 35 and 52]) and the condition numbers are the diagonal elements of the overlap matrix defined by the left and the right eigenvectors for non-Hermitian matrices. The coupling between the eigenvalue process and the eigenvalue-overlap process is essential to describe the non-Hermitian matrix-valued BM [12, 13, 14, 15, 18, 21, 29].
(iv)   It is an important problem to study the present models by analyzing the right- and the left-eigenvectors [32]. Moreover, we shall study the non-Hermitian matrix-valued stochastic processes starting from nonnormal matrices with defective eigenvalues from the viewpoint of the coupling between the eigenvalue proceses and the pseudospectum processes. We expect that if we impose special symmetries to the non-Hermitian matrix-valued processes, e.g., by putting all entries are real or quaternions, both of the eigenvalue distributions and pseudospectra become properly structured [39, Section 50].
Recently non-Hermitian quantum physics has been extensively studied (see [3] and the references therein). We expect that the present mathematical study of eigenvalue and pseudospectrum processes will be related to the pseudospectrum approaches to non-Hermitian systems discussed in physics [28, 33].
Acknowledgements   This study was carried out under the Open-Type Professional Development Program of the Institute of Statistical Mathematics, Tokyo (2023-ISMHRD-7010) organized by Takaaki Shimura, Satoshi Kuriki, Benoît Collins, Noriyoshi Sakuma, and Yuki Ueda. The present authors would like to thank J. Garza-Vargas for useful comments on the present work. Part of the present study was done during the stay of the authors at Institute for Mathematical Sciences, National University of Singapore. The authors thank Akira Sakai and Rongfeng Sun for organizing the three-week fruitful program in December 2023. This work was also supported by the Research Institute for Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University. MS was supported by JSPS KAKENHI Grant Numbers JP19K03674, JP21H04432, JP22H05105, JP23K25774, and JP24K06888. TS was supported by JSPS KAKENHI Grant Numbers JP20K20884, JP22H05105 and JP23K25774. TS was also supported in part by JSPS KAKENHI Grant Numbers JP21H04432 and JP24KK0060.
Appendix A A Proof of Lemma 3.6
For $n\in\mathbb{N}$, let
$p_{1}(n,m):=\left[\frac{n}{m}\right],\quad p_{2}(n,m):=\left[\frac{n}{m+1}\right].$
Here we prove the following statement. If we replace $n$ by $n-1$, then Lemma 3.6 in Section 3.2 is obtained.
Lemma A.1
Let $I_{n}:=[\lceil\sqrt{n}\,\rceil,n]\cap\mathbb{N}$ and $T_{n}:=\{[n/k];k=1,2,\dots,n\}$. Then
$p_{1}(n,m)-p_{2}(n,m)=\begin{cases}1,\quad\mbox{if $m\in I_{n}\cap T_{n}$},\cr 0,\quad\mbox{if $m\in I_{n}\setminus T_{n}$}.\end{cases}$
Remark A.1   Let $T_{n}^{\prime}:=\{[n/k];k=1,2,\dots,[\sqrt{n}\,]-1\}$ and $T_{n}^{\prime\prime}:=\{[n/k];k=1,2,\dots,[\sqrt{n}\,]\}$. Then $T_{n}^{\prime}\subset I_{n}\cap T_{n}\subset T_{n}^{\prime\prime}$ for any $n$. Indeed,
$n\geq([\sqrt{n}\,]-1)(\sqrt{n}+1)\Longrightarrow\frac{n}{[\sqrt{n}\,]-1}\geq\sqrt{n}+1\Longrightarrow\left[\frac{n}{[\sqrt{n}\,]-1}\right]\geq\lceil\sqrt{n}\,\rceil$
and
$\left[\frac{n}{k}\right]\geq\lceil\sqrt{n}\,\rceil\Longrightarrow\frac{n}{k}\geq\sqrt{n}\Longrightarrow\sqrt{n}\geq k.$
Proof of Lemma A.1   From Remark A.1 above, it is sufficient to show that
  • (i)
    $p_{1}(n,m)-p_{2}(n,m)\in\{0,1\}$ if $m\in I_{n}$.
  • (ii)
    If $p_{1}(n,m)-p_{2}(n,m)=1$, then $m\in T_{n}$.
  • (iii)
    If $m\in T_{n}^{\prime\prime}$, then $p_{1}(n,m)-p_{2}(n,m)=1$.
(i) First we remark that $p_{1}(n,m)-p_{2}(n,m)\in\{0,1\}$ if $m\in I_{n}$. Indeed, since $m\geq\lceil\sqrt{n}\,\rceil\geq\sqrt{n}$, we see that
$20\leq p_{1}(n,m)-p_{2}(n,m)<\frac{n}{m(m+1)}+1\leq\frac{n}{\sqrt{n}(\sqrt{n}+1)}+1<$ 2 .
(ii) Suppose $p_{1}(n,m)-p_{2}(n,m)=1$, i.e., $p_{1}(n,m)=k$ and $p_{2}(n,m)=k-1$ for some $k=1,2,\dots,n$. Then, the following conditions hold.
$n$ $=km+r\quad(\exists r\in\{0,1,\dots,m-1\})$
$n$ $=(k-1)(m+1)+r^{\prime}\quad(\exists r^{\prime}\in\{0,1,\dots,m\}),$
which implies that
$km\leq n\leq km+m-1=(k+1)m-1$
$(k-1)(m+1)\leq n\leq(k-1)(m+1)+m=km+k-1,$
and hence
$km\leq n\leq km+k-1\Longrightarrow m=[n/k].$
Therefore, if $p_{1}(n,m)=k$ and $p_{2}(n,m)=k-1$, then $m=[n/k]$. In other words,
$\{m\in\{1,2,\dots,n\}:p_{1}(n,m)-p_{2}(n,m)=1\}\subset T_{n}.$
(iii) Suppose $m\in T_{n}^{\prime\prime}$, i.e, $m=[n/k]$ for some $k=1,2,\dots,[\sqrt{n}\,]$. We see that
$\frac{n-k+1}{k}\leq\left[\frac{n}{k}\right]\leq\frac{n}{k}$
and thus
$k\leq\frac{n}{[\frac{n}{k}]}\leq\frac{nk}{n-k+1}=k+\frac{k(k-1)}{n-k+1}.$
If $1\leq k\leq[\sqrt{n}\,]$, then
$\frac{k(k-1)}{n-k+1}\leq\frac{n-\sqrt{n}}{n-\sqrt{n}+1}<1,$
which implies
$\left[\frac{n}{m}\right]=\left[\frac{n}{[\frac{n}{k}]}\right]=k.$
Similarly, we see that
$\frac{n+1}{k}\leq\left[\frac{n}{k}\right]+1\leq\frac{n+k}{k}$
and thus
$\frac{nk}{n+k}\leq\frac{n}{[\frac{n}{k}]+1}\leq\frac{nk}{n+1}<k.$
If $1\leq k\leq[\sqrt{n}\,]$, then
$\frac{nk}{n+k}=k-1+\left(1-\frac{k^{2}}{n+k}\right)\geq k-1,$
which implies
$\left[\frac{n}{m+1}\right]=\left[\frac{n}{[\frac{n}{k}]+1}\right]=k-1.$
Therefore, if $1\leq k\leq[\sqrt{n}\,]$, then
$p_{1}(n,m)-p_{2}(n,m)=k-(k-1)=1.$